OpenStudy (anonymous):

simplify. 4√10mn • 3√5m

OpenStudy (paxpolaris):

are the m's and n inside the radical

OpenStudy (paxpolaris):

?

OpenStudy (anonymous):

yes

OpenStudy (paxpolaris):

\[\Large 4\sqrt{10mn} \times 3\sqrt{5m}\]\[\Large = 12 \sqrt{2\cdot5\cdot5\cdot m^2n}\]

OpenStudy (paxpolaris):

\[\Large = 12 \cdot 5 m \sqrt {2n}=\LARGE 60m \sqrt {2n}\]